### v1

@@ 5.4 Utilizing Machine Learning
We now have the features dataset and UOCU (as labels) ready to run a machine learning operation. The goal of this part is to discover if there is a relationship between the metrics of notebooks and their CU. It is also possible to provide a model that can estimate the level of understandability of a notebook. As the dataset features are all numbers and are not that complex, we employed four machine learning models to discover the relationship. We employed Random Forest, CatBoost, Decision Tree, and XGBoost models for the classification of notebooks into binary classes. CatBoost is a gradient boosting algorithm that handles categorical features efficiently, while XGBoost is known for its scalability, speed, and ability to handle missing data. Decision trees are interpretable and can capture complex relationships within the data. Finally, Random Forest combines multiple decision trees to improve accuracy and reduce overfitting, making it particularly useful when dealing with high-dimensional data and noisy features. By experimenting with these diverse algorithms, we aimed to find the best-performing model for our specific task of predicting notebook CU. Additionally, their simplicity allowed us to concentrate on the analysis of the features rather than intricate model architectures or complex hyperparameter tuning. The results of the evaluation metrics and the accuracy of these algorithms are presented in Section 6.2.3. The Random Forest algorithm ultimately stood out with an accuracy of 89% and F1-score of 88%. We conducted a grid search on a subset of the data to explore different hyperparameter configurations. We found that the default settings for these algorithms yielded competitive results while significantly reducing the computational overhead.
@@ 6.1.1 Sampling
To streamline and optimize the study, a filter was applied to eliminate notebooks. After extensive experimentation, we observed that UOCU typically does not yield favorable scores for notebooks with fewer than 500 views. This phenomenon likely arises from the potential for an increase in comments as views accumulate. Therefore, we removed them from the 542,051 notebooks in the DistilKaggle dataset, resulting in a reduced count of 132,723 notebooks. Based on calculations similar to those outlined by Zhu et al. zhu2021restoring , we determined that a sample size of 1,050 notebooks is required to attain a confidence level of 99% with a margin of error of ±4% sample2024 . Consequently, these 1,050 notebooks were extracted and excluded from the training set and preserved to establish ground truth.
6.1.2 Human Labeling
We designed an experiment to assess the performance of the CU scores introduced in this paper. We conducted a human labeling task in which each annotator categorized notebooks into “good” and “normal” CU categories, following the definition provided by Xia et al. in Section 1.
Experts were assigned to annotate the sample dataset outlined in Section 6.1.1. Each notebook was evaluated by three experts and classified as either ‘NCU’ (normal code understandability) or ‘GCU’ (good code understandability).
@@ 6.1.2 Human Labeling
To establish a ground truth dataset, we followed the labeling process methodology employed by Guzman et al. guzman2015ensemble . Initially, we developed a set of instructions containing definitions for each class and several examples showcasing various characteristics such as line count, comments, markdown, different coding styles, display outputs, etc. This guideline assisted annotators in understanding the classification process and helped minimize disagreements. Subsequently, a random selection of 1,050 notebooks detailed in Section 6.1.1 was made.
We invited 55 experts to participate, and eventually received responses from 42 of them. We used convenience sampling to select these experts. They were selected for inclusion in the sample because they were the easiest for us to access. Through the information gathered from these participants, we determined an average experience level of 5 years working with Python in scientific or industrial contexts. The notebook samples were systematically selected by the authors and randomly assigned to experts, each assessed by three annotators to ensure comprehensive evaluation and reliability of results. To ease the review process and conceal metadata like user comments, upvotes, and views, the notebooks were placed on nbviewer.com. The annotators conducted a manual analysis of the notebooks and assigned the appropriate labels. Subsequently, we addressed any disagreements in the reviews and conducted a meeting to reach final decisions, particularly for annotators with unlabeled notebooks.
In order to ensure the reliability and validity of our questionnaire, we employed kappa coefficients as a measure of inter-rater agreement. Given the involvement of three raters in our task, Cohen’s kappa was not applicable due to its restriction to two raters. Therefore, we utilized Fleiss’ kappa, which accommodates more than two raters. This approach allowed us to assess the consistency among multiple raters and provide robust assurance of the questionnaire’s quality. More details about Fleiss’ kappa are described in Section 6.1.4.
This comprehensive evaluation yielded a Fleiss’ kappa coefficient of 82%, providing almost perfect agreement of the questionnaire’s quality and reliability landis1977application . The labeled notebooks comprised 566 (54%) labeled as ‘NCU’ and 484 (46%) labeled as ‘GCU’.
@@ 6.1.3 Baselines
For our dataset, we needed to find models and criteria suitable for the specific task of CU classification of Jupyter notebooks. To accomplish this, we selected three different methods:
@@
CodeBERT has been pre-trained in both programming language (PL) and natural language (NL). It learns representations that are useful for various NL-PL applications such as natural language code search and code documentation generation. PLBART (Pretrained LAnguage- and Vision-based BART) is a state-of-the-art deep model that extends BART to incorporate both language and vision inputs, enabling it to generate high-quality text in multiple languages while understanding visual information. We employed the Hugging Face Transformers library for our work huggingface2019transformers . We did not have access to labeled data for fine-tuning these transformers. We only had 1,050 notebooks as our ground truth, which were used as our final test dataset to validate the performance of our methods. To address the scarcity of training data, we employed a straightforward approach of ranking notebooks based on their upvotes, considering it as a rudimentary metric for assessing CU. This approach aligns with previous studies, which often consider highly-voted notebooks as a proxy for well-documented notebooks Wang2021WhatMakes ; mondal2023cell2doc , or utilize the number of stars and forks as proxies for the quality of data science GitHub projects morakot2023mining ).
@@
Our analysis, as depicted in Figure 3, revealed a trend wherein increasing the sample size for fine-tuning baseline models led to improved performance, up to a threshold of 4,000 samples. Beyond this threshold, we observed no significant improvements in model training performance. Based on these findings, we concluded that a total sample size of 4,000 notebooks would be optimal for our study, comprising 2,000 notebooks with the lowest upvote scores, referred to as ’NCU,’ and an additional 2,000 notebooks with the highest scores, referred to as ’GCU.’
@@
Regarding training parameters, we employed accuracy metric to assess model performance. We configured the “evaluation_strategy” as “epochs” and conducted training over 10 epochs. We restricted the number of epochs to 10, as a result of an observation that the model reached convergence on the input data within 10 epochs. We also set arguments regarding whether or not to run training on the training set, evaluation on the validation set, and predictions on the test set all to “True”. Notably, our code was executed on a system with a GPU, as detailed in Section 6.1.
@@ 6.1.4 Evaluation Metrics
We utilize Accuracy, Precision, Recall, and F1-score to evaluate our classifiers. Accuracy calculates the percentage of all notebooks that are predicted correctly (as either GCU or NCU) out of all the existing notebooks in our ground truth. Precision calculates the percentage of notebooks that are predicted correctly as GCU among all notebooks predicted as GCU (whether correct or not). Recall calculates the percentage of notebooks that are predicted correctly as GCU among all notebooks that are labeled as GCU in our ground truth. The F1-score assesses the harmonic mean of precision and recall. Since both GCU and NCU are equally important, the F1-score serves as an appropriate evaluation metric that demands our consideration.
@@ 6.2.1 Evaluation of the CU Criteria
    The hybrid approach: In this approach, we remove outlier data, scale the scores of approach 1 and 4, and then consider their sum as a criterion.
@@
Answer: Based on the results we achieved, it is possible to predict the CU scores of the notebooks on the Kaggle to a great extent based on their metadata. With a hybrid approach based on the UOCU score and upvoted notebooks, we achieved an F1-score of 86% and accuracy of 87% in predicting CU category, which represents a significant improvement over other baselines.
6.2.2 Evaluating the Predictive Model
@@
Following the procedures outlined in Section 5.2 for calculating notebook CU scores in binary classification mode, in order to enhance the accuracy of classifying notebooks into GCU and NCU categories, we omitted the second and third quartiles (the middle two quartiles). As a result, a total of 66,360 (out of 132,723) notebooks remained in the dataset, comprising 33,180 notebooks labeled as NCU and 33,180 notebooks labeled as GCU. In preparation for training, we divided these notebooks into three distinct sets: the training set, the evaluation set, and the test set, with ratios of 70%, 20%, and 10% respectively. In the training phase, we employed the evaluation set to mitigate overfitting and utilized the test set to optimize the hyperparameters of our candidate machine learning models. Ultimately, to determine the optimal model, we assessed the trained models using our dataset of 1,050 ground truth notebooks as introduced in Section 6.1.2. All numerical results presented for machine learning algorithm outcomes in binary classification are based on this ground truth dataset.
Before delving into the examination of the accuracy levels in predicting binary classes using machine learning algorithms, it’s worth exploring the accuracy levels of these algorithms across two different modes: four-level classification, and three-level classification.
In the quaternary classification, we reintroduced the two sets of notebooks that were excluded during binary classification as two additional classes. The prediction results for this classification are outlined in Table 5.
@@
In the context of ternary classification, we evenly divided the data into three equal parts based on their CU scores. You can find the evaluation metric results in Table 6.
@@
Now that we’ve assessed the prediction accuracy results for the three scenarios mentioned earlier, we can shift our focus back to evaluating the prediction accuracy specifically for binary classification.
The prediction results of the evaluation for our final dataset using four machine learning algorithms are presented in Table 7.
@@
As shown in Table 7, Random Forest produces the best results with an F1-score of 88% and an Accuracy score of 89%. We also utilize the AUC-ROC criterion, as introduced in Section Section 6.1.4, to provide additional validation for our experiments. Notably, this criterion yields a value of 94% for the Random Forest algorithm, reinforcing our confidence in its selection.
@@
RQ2. Is it feasible to develop a model capable of achieving satisfactory performance in assessing the code understandability of notebooks?
Answer: We developed a predictive model using the Random Forest machine learning algorithm to evaluate CU in notebooks through binary classification. This endeavor yielded a commendable accuracy of 89% along with an F1-score of 88%, showcasing significant improvements in accuracy and F1-score compared to the baseline.
@@ 6.2.3 Finding Features Effective on CU
In this study, we aimed to investigate the factors that impact CU in a dataset of notebooks. To achieve this, we applied the SHAP framework to assess the importance of different features, as discussed in Section 6.1.5. Our results are presented in Figure 4, which illustrates a Beeswarm summary box plot. This plot offers a condensed overview of how the most significant notebook features impact the model’s output.
@@
Perhaps we are looking for a model capable of predicting the CU score using static metrics extracted from a notebook, regardless of the developer’s skill level. In this context, assigning a value to the PT metric might not be straightforward. In a separate experiment, we omitted this metric and attained an accuracy and F1-score of 80% using the Random Forest algorithm.
@@
We continued to investigate the impact of each group or subgroup of code metrics on code comprehensibility. The results of this investigation are provided in Table 9. As evident, notebook-based metrics are more effective on their own compared to script-based metrics. Moreover, within the category of script-based metrics, basic and halstead metrics demonstrate greater effectiveness.
@@
RQ3. Which notebook metrics can affect CU, and to what extent can each metric influence code understandability?

### v2

@@ 5.4 Utilizing Machine Learning
We now have the features dataset and UOCU (as labels) ready to run a machine learning operation. The goal of this part is to discover if there is a relationship between the metrics of notebooks and their CU. It is also possible to provide a model that can estimate the level of understandability of a notebook. As the dataset features are all numbers and are not that complex, we employed four machine learning models to discover the relationship. We employed Random Forest, CatBoost, Decision Tree, and XGBoost models for the classification of notebooks into binary classes. CatBoost was included because of its native support for categorical features without the need for extensive preprocessing, making it particularly suited to our dataset’s structure. Additionally, it has demonstrated strong performance and reduced computational costs compared to other boosting methods. While XGBoost is known for its efficiency, scalability, and exceptional performance in structured/tabular data problems. It incorporates advanced features like regularization and tree pruning, making it highly effective in minimizing overfitting. Decision trees are interpretable and can capture complex relationships within the data. Finally, Random Forest was included due to its robustness against overfitting, strong performance with high-dimensional datasets, and ability to handle both classification and regression tasks effectively. Its ensemble nature leverages the power of multiple decision trees to improve accuracy and reduce variance. By experimenting with these diverse algorithms, we aimed to find the best-performing model for our specific task of notebook CU classification. Additionally, their simplicity allowed us to concentrate on the analysis of the features rather than intricate model architectures or complex hyperparameter tuning. The results of the evaluation metrics and the accuracy of these algorithms are presented in Section 6.2.3. The Random Forest algorithm ultimately stood out with an accuracy of 89% and F1-score of 88%. We conducted a grid search on a subset of the data to explore different hyperparameter configurations. We found that the default settings for these algorithms yielded competitive results while significantly reducing the computational overhead.
@@ 6.1.1 Sampling
To streamline and optimize the study, a filter was applied to eliminate notebooks. In our initial exploration of the DistilKaggle dataset, we observed a wide range of notebook view counts, from just a few views up to thousands. We systematically evaluated the performance of our UOCU criterion on notebooks with varying view counts. Specifically, we randomly sampled notebooks from the dataset and divided them into bins based on their view counts (e.g., 0-100 views, 100-500 views, 500-1000 views, and so on). We then calculated UOCU for each bin and found that it results in zero values for most of them. We then manually inspected these bins to further investigate the findings. Our findings consistently showed that the UOCU approach struggled to produce scores for notebooks with fewer than 500 views. We hypothesize that this is due to the relatively low number of comments and user interactions associated with these notebooks. These notebooks have the potential to attract many more comments over time, but for now, we decided to apply a 500-view filter to the dataset. This filtering step resulted in a reduction in the dataset size, from 542,051 to 132,723 notebooks.
Based on sampling calculations similar to those outlined by Zhu et al. zhu2021restoring , we determined that a sample size of 1,050 notebooks is required to attain a confidence level of 99% with a margin of error of ±4%. Specifically, we applied the sample size formula provided by Creative Research Systems sample2024 , which aligns with the methodological approach used by Zhu et al., while adjusting for the parameters of our study. Consequently, these 1,050 notebooks were extracted and excluded from the training set and preserved to establish ground truth.
6.1.2 Ground Truth
We designed a human labeling task to classify computational notebooks into two categories: NCU and GCU. This task evaluates the CU scores introduced in our study, following the annotation framework outlined in Appendix D. Specifically, annotators assessed aspects such as logical structure, descriptive code and documentation, and error handling—key factors influencing code comprehensibility. These characteristics were derived from prior research to ensure alignment with established standards of code understandability rule2018exploration ; pimentel2019large ; perkel2018jupyter .
@@ 6.1.2 Ground Truth
To establish a ground truth dataset, we adopted the labeling methodology of Guzman et al.guzman2015ensemble on the dataset detailed inSection 6.1.1. Initially, we provided annotators with a set of instructions defining each class, along with examples illustrating various characteristics. This guideline helped standardize the classification process and minimize disagreements.
We invited 55 experts to participate in the labeling task, receiving responses from 42. These experts had an average of five years of experience working with Python in scientific or industrial contexts. Notably, over 90% were familiar with Jupyter notebooks, meeting the criteria necessary for accurate labeling.
Notebook samples were randomly assigned to experts, with each notebook assessed by three annotators to ensure reliability. To facilitate review while concealing metadata such as user comments, upvotes, and views, the notebooks were hosted on nbviewer.org, an online tool for viewing Jupyter notebooks in a browser without requiring a local installation. Annotators conducted manual analyses and assigned labels accordingly. Disagreements were addressed through discussions, and a final consensus was reached, particularly for cases with missing labels.
To measure inter-rater agreement and ensure labeling reliability, we employed Fleiss’ kappa, which accommodates multiple raters. This was necessary as Cohen’s kappa applies only to two raters. More details on Fleiss’ kappa are provided in Section 6.1.4. Our evaluation yielded a Fleiss’ kappa coefficient of 82%, indicating almost perfect agreement and confirming the reliability of our labeling process landis1977application . The final dataset consisted of 566 notebooks (54%) labeled as NCU and 484 (46%) as GCU. We used these 1,050 labeled notebooks as ground truth to manually validate our UOCU criterion and other baselines, with results reported in Section 6.2.1.
@@ 6.1.3 Baselines
For our dataset, we needed to find models and criteria suitable for the specific task of CU classification of Jupyter notebooks.
As outlined in Section 8, our study aligns with prior work that measures CU based on human opinions, including subjective evaluations, surveys, and community-based metrics. However, existing approaches have notable limitations:
Given these limitations, we chose upvotes as a baseline due to their prevalence in repository-based evaluations of CU.
To accomplish this, we selected three different methods:
@@
CodeBERT has been pre-trained in both programming language (PL) and natural language (NL). It learns representations that are useful for various NL-PL applications such as natural language code search and code documentation generation. PLBART (Pretrained LAnguage- and Vision-based BART) is a state-of-the-art deep learning model that extends BART to incorporate both language and vision inputs, enabling it to generate high-quality text in multiple languages while understanding visual information. We employed the Hugging Face Transformers library for our work huggingface2019transformers . We did not have access to labeled data for fine-tuning these transformers. We only had 1,050 notebooks as our ground truth, which were used as our final test dataset to validate the performance of our methods. To address the scarcity of training data, we employed a straightforward approach of ranking notebooks based on their upvotes, considering it as a rudimentary metric for assessing CU. This approach aligns with previous studies, which often consider highly-voted notebooks as a proxy for well-documented notebooks Wang2021WhatMakes ; mondal2023cell2doc , or utilize the number of stars and forks as proxies for the quality of data science GitHub projects morakot2023mining ).
@@
Our analysis, as depicted in Figure 3, revealed a trend wherein increasing the sample size for fine-tuning baseline models led to improved performance, up to a threshold of 4,000 samples. Beyond this threshold, we observed no significant improvements in model training performance. Based on these findings, we concluded that a total sample size of 4,000 notebooks would be optimal for our study, comprising 2,000 notebooks with the lowest upvote scores, referred to as ‘NCU’, and an additional 2,000 notebooks with the highest scores, referred to as ‘GCU’.
@@
Regarding training parameters, we employed accuracy metric to assess model performance. We configured the “evaluation_strategy” as “epochs” and conducted training over 10 epochs. We restricted the number of epochs to 10, as a result of an observation that the model reached convergence on the input data within 10 epochs. We also set arguments regarding whether or not to run training on the training set, evaluation on the validation set, and classification on the test set all to “True”. Notably, our code was executed on a system with a GPU, as detailed in Section 6.1.
@@ 6.1.4 Evaluation Metrics
We utilize Accuracy, Precision, Recall, and F1-score to evaluate our classifiers. Accuracy calculates the percentage of all notebooks that are predicted correctly (as either GCU or NCU) out of all the existing notebooks in our ground truth. It is calculated as Equation (2) where True Positives are the correctly predicted GCUs, and True Negatives are the correctly predicted NCUs. Precision, in the context of GCUs as positive labels, measures the proportion of correctly predicted GCU notebooks out of all notebooks predicted as GCUs. It is calculated as Equation (3) where False Positives are the NCUs that were incorrectly predicted as GCUs. Recall, also known as Sensitivity or True Positive Rate, is the proportion of correctly predicted GCU notebooks out of all actual GCU notebooks in the dataset. It is calculated as Equation (4) where False Negatives are the GCUs that were incorrectly predicted as NCUs. The F1-score is the harmonic mean of precision and recall. It provides a single metric that balances the trade-off between precision and recall, especially useful when dealing with class imbalance. It is calculated as Equation (5).
@@
$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ (2)
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (3)
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (4)
[…]
@@ 6.2.1 Evaluation of the CU Criteria
    The hybrid approach: In this approach, we address outliers, normalize the scores from Approach 1 and Approach 4, and then consider their sum as a new evaluation criterion. Outliers were identified based on the number of upvotes and the UOCU score. Specifically, approximately 1% of the data with both low upvotes and low UOCU scores were removed. To normalize the scores from Approach 1 and Approach 4, we applied Min-Max normalization  Aksu2019normalization using the formula in Equation (6):
    $A^{\prime}=\frac{A-\min(A)}{\max(A)-\min(A)}$ (6)
    where $A$ represents the original score, and $A^{\prime}$ is the normalized score. After normalization, the scores from both approaches were aggregated to create a hybrid score.
@@
Answer: Based on the results we achieved, it is possible to predict the CU scores of the notebooks on Kaggle to a great extent based on their metadata. With a hybrid approach based on the UOCU score and upvoted notebooks, we achieved an F1-score of 86% and accuracy of 87% in predicting CU category, which represents a significant improvement over other baselines.
6.2.2 Evaluating the Classifier
@@
Based on Section 6.2.1, we found the best strategy for evaluation of CU, the hybrid approach. As mentioned in Section 5.2, we had calculated the hybrid CU scores in binary classification mode, and omitted the second and third quartiles (the middle two quartiles). As a result, a total of 66,360 (out of $\sim 132,720$) notebooks remained in the dataset, comprising 33,180 notebooks labeled as NCU and 33,180 notebooks labeled as GCU. The range of scores defining each quartile and the corresponding number of notebooks are as shown in Table 6.
Table 6: Range of Hybrid CU Scores and The Number of Notebooks in Each Quartile
Quartile CU Score Range Number of Notebooks
Q1 0.01 – 0.06 33,180
Q2 0.06 – 0.08 33,180
Q3 0.08 – 0.11 33,180
Q4 0.11 – 1.84 33,180
In preparation for training, we divided these notebooks into three distinct sets: the training set, the evaluation set, and the test set, with ratios of 70%, 20%, and 10% respectively. In the training phase, we employed the evaluation set to mitigate overfitting and utilized the test set to optimize the hyperparameters of our candidate machine learning models. Ultimately, to determine the optimal model, we assessed the trained models using our dataset of 1,050 ground truth notebooks as introduced in Section 6.1.2. All numerical results presented for machine learning algorithm outcomes in binary classification are based on this ground truth dataset.
Before examining the accuracy of machine learning algorithms in binary classification, it is useful to first evaluate their accuracy across two different modes: four-level classification and three-level classification.
In our analysis, we employed three classification modes to evaluate the CU of notebooks:
  • Binary Classification: This mode categorizes notebooks into two classes: Good Code Understandability (GCU) and Not Code Understandable (NCU).
  • Ternary Classification: In this mode, notebooks are divided into three equal parts based on their CU scores, allowing for a more nuanced understanding of their quality.
  • Quaternary Classification: This classification reintroduces two additional classes that were excluded in the binary classification, resulting in four distinct categories for evaluation.
Table 7 presents the classification results for the quaternary classification of notebooks, detailing the accuracy and F1-score achieved by our machine learning models across the four defined categories.
@@
Table 8 presents the evaluation metrics for ternary classification, illustrating the performance of the models in classifying notebooks into three equal classes.
@@
Now that we’ve assessed the classification accuracy results for the three scenarios mentioned earlier, we can shift our focus back to evaluating the classification accuracy specifically for binary classification.
The classification results of the evaluation for our final dataset using four machine learning algorithms are presented in Table 9.
@@
As shown in Table 9, Random Forest produces the best results with an F1-score of 88% and an Accuracy score of 89%. We also utilize the AUC-ROC criterion, as introduced in Section 6.1.4, to provide additional validation for our experiments. Notably, this criterion yields a value of 94% for the Random Forest algorithm, reinforcing our confidence in its selection.
@@
RQ2. RQ2: Can a model be developed that effectively classify code understandability of notebooks based on the evaluation metrics established in this study?
Answer: We developed a Classifier using the Random Forest machine learning algorithm to evaluate CU in notebooks through binary classification. This endeavor yielded a commendable accuracy of 89% along with an F1-score of 88%, showcasing significant improvements in accuracy and F1-score compared to the baseline.
@@ 6.2.3 Finding Features Effective on CU
In our study, we aimed to investigate the factors that impact CU in a dataset of notebooks. To achieve this, we applied the SHAP framework to assess the importance of different features, as discussed in Section 6.1.5. Our results are presented in Figure 4, which illustrates a Beeswarm summary box plot. This plot offers a condensed overview of how the most significant notebook features impact the model’s output.
@@
Perhaps we are looking for a model capable of classification using static metrics extracted from a notebook, regardless of the developer’s skill level. In this context, assigning a value to the PT metric might not be straightforward. In a separate experiment, we omitted this metric and attained an accuracy and F1-score of 80% using the Random Forest algorithm.
@@
We continued to investigate the impact of each group or subgroup of code metrics on code comprehensibility. The results of this investigation are provided in Table 11. This analysis aimed to determine which algorithm can best assess code understandability with the highest accuracy for each group of code metrics. The findings reveal that notebook-based metrics are more effective on their own compared to script-based metrics. Furthermore, within the category of script-based metrics, basic and Halstead metrics exhibit greater effectiveness.
@@
RQ3. RQ3: Which notebook metrics are associated with code understandability, and how significantly do they influence the classification of code understandability?
@@
7 Discussion
This section summarizes the results, compares them with existing findings, discusses potential applications, and highlights the novelty of our approach.
Findings: The study consists of two phases; The first phase introduced a method to assess CU in Jupyter notebooks using Kaggle metadata and user feedback. Five approaches were tested, including fine-tuned CodeBERT, PLBART, upvote-based models, the UOCU approach, and a hybrid model combining the latter two. The hybrid model performed best, achieving 87% accuracy. The second phase analyzed which notebook metrics impact CU the most. Machine learning models were applied to CU scores, with SHAP analysis identifying key factors. Notebooks from proficient users were more understandable, while excessive blank lines reduced clarity. Metrics like visualization outputs, executed cells, top-level headers, and markdown lines positively influenced CU, highlighting the importance of clean code, documentation, and visual aids. Additionally, notebook-based metrics (markdown quality, execution clarity, visualizations) had a stronger impact on CU than traditional script-based metrics, emphasizing the unique, interactive nature of computational notebooks.
Comparison to Existing Work: Our findings underscore the performance tier (PT) as a significant factor influencing CU. This aligns with prior studies that suggest developer expertise impacts code quality, understandability or refactoring. For instance, Alomar et al. alomar2021behind show that experienced developers are more inclined to refactor, improving code quality and readability. Similarly, Scalabrino et al.scalabrino2019automatically reveal that CU is often more affected by the developers’ proficiency than by specific code snippets.
Additionally, the positive effect of metrics such as Header 1 and Lines of Markdown Cells on CU underscores the importance of documentation, particularly the use of markdown cells, in enhancing notebook CU. Prior studies support this finding, highlighting the role of documentation in improving readability and understandability. For example, Börstler et al. borstler2023developers examine how practitioners perceive and discuss various aspects of code quality, including understandability. Documentation is identified as a key factor influencing CU, with developers frequently associating well-documented code with higher quality. Furthermore, a considerable number of previous studies have highlighted the positive impact of markdown on code understandability. However, unlike our study, the extent of this effect has not been quantitatively evaluated on a large dataset pimentel2019large ; venkatesh2021automated ; dong2021splitting .
Also, our study provides novel insights into how the number of executed cells and the visualization data types in Cell Output specifically affect CU in computational notebooks. These metrics, which are unique to the notebook environment, relate directly to the outputs generated by code cells. Our findings indicate that when the outputs are limited or when there is a scarcity of visual elements—such as images, plots, and diagrams—the understandability of the notebook decreases. This underscores the essential role of visualizations in enhancing code understandability, as they provide intuitive context that aids in interpreting results and understanding the flow of analysis.
Novelty and Contributions: Our study provides a novel perspective by leveraging user feedback as a metric for measuring CU, an approach not previously explored in code understandability research. Unlike traditional methods that primarily focus on code structure or static analysis, our study integrates user opinions as a core component in assessing the CU of Jupyter notebooks—a widely-used format in data science and machine learning that has not been extensively explored in CU research. What distinguishes our approach is the analysis of comments and feedback directly shared by users within notebook repositories, adding an additional layer of insight. This inclusion captures a broad spectrum of real-world perceptions and interactions from a large user base, which is often absent in methods that rely solely on structural metrics or small-scale, survey-based feedback.
Furthermore, after determining the most effective method for measuring CU (the hybrid approach combining user opinions and notebook upvotes), we employed machine learning techniques to identify the notebook metrics that most significantly influence CU. By utilizing multiple machine learning algorithms, we presented, for the first time, a model capable of assessing notebook CU levels. This enabled us to identify the most impactful metrics, such as performance tier, markdown quality, and visualization elements, providing actionable insights for improving notebook comprehensibility.
Implications and Potential Applications: This section outlines the applications of our approach and the generalizability of our findings. By leveraging user feedback as a criterion for CU, this study highlights the potential for applying similar methods across various code repositories, platforms, and programming languages. Integrating user insights into CU assessments introduces a valuable context-sensitive criterion, creating a more holistic, user-centered understanding of code quality.
Our findings offer significant implications for notebook authors and tool developers aiming to enhance CU in Jupyter notebooks. Recognizing the influence of performance tier as a key factor suggests that resources like tutorials, training modules, and mentorship programs could help novice users develop coding expertise, thereby improving overall notebook quality. Additionally, notebook environments could introduce expertise-driven suggestions—such as reminders for more documentation or alerts about excessive blank lines—to support better readability, as our study found these aspects can detract from clarity. Furthermore, the substantial impact of notebook-specific metrics, including markdown cells, headers, and visualizations, underscores the need for tools that encourage well-organized, interactive analyses. Enhancing applications like JupyterLab with extensions for visualization and markdown formatting can guide users in structuring their notebooks more effectively. This is particularly valuable in educational settings, where notebooks are widely used for teaching programming and data science. By promoting the use of rich visual content, such as plots, images, and interactive components, developers can help users create notebooks that are both functional and visually intuitive, ultimately facilitating a deeper understanding of complex analyses.